OpenStudy (anonymous):

like if a man has 500 ft of fencing what is the largest area that can be made if the fence is totally enclosed

OpenStudy (radar):

I could just say a square with 125 ft. per side is the largest area he can get. but I bet you would want proof.

OpenStudy (radar):

Here is the proof: Let x=one side, so the opposite is also x Amount of fence left is 500-2x so that will be divided between the other two opposite sides. 250-x = other side. A= (x)(250-x)=-x^2 + 250x Now we differentiate and set to equal zero -2x+250=0 (max/min) -2x=-250 x=125

OpenStudy (anonymous):

actually thts correct i just figure it out its formulating the equations tht i have trouble with derivatives r no problem

OpenStudy (anonymous):

thank u thats the easy one but what about when there are four enclosed areas of equal size? i know the derivatives but finding the equation is urksome

OpenStudy (radar):

When it gets like that, i need a diagram lol

OpenStudy (anonymous):

oh ok